Calculate the amplitude, period, phase shift, and use the information to sketch one full cycle of the graph of the equation f(t) = 0.2 cos (pi/12 t - 7pi/12), which is used in predicting the height of ocean tidal components.

Answer by stanbon(75887)

Calculate the amplitude, period, phase shift, and use the information to sketch one full cycle of the graph of the equation f(t) = 0.2 cos (pi/12 t - 7pi/12), which is used in predicting the height of ocean tidal components.
----
If y = a*sin(bx+c)+d
------
amp = |a| = 0.2
---
period = (2pi)/b = (2pi)/(pi/12) = 24
---
phase shift = -c/b = -(-7pi/12)/(pi/12) = 7
==================
Cheers,
Stan H.
----